OpenStudy (anonymous):

10 mLof 3.0M HCl to grams?

OpenStudy (anonymous):

you gotta remember those godly formulas lol! Molarity=mole/volume and Mole=mass/mr first to do it, find the mole of HCl >> since we know the molarity and the volume, mole=molarity x volume. means, 3 mole dm^-3 x 0.01 dm^3 (convert 10 mL to L or dm^3). the result is 0.03 mole second, put the mole to this formula "mole=mass/mr". 0.03 mole=x (we dont know what's the mass, just put x there) over Mr which is 1+35.5=36.5, resulted to 1.095 gr

OpenStudy (anonymous):

hoping it's right though. do you understand the whole point?

OpenStudy (anonymous):

Yes. Thanks, I just hate stoichiometry.

OpenStudy (anonymous):

sure ;) glad to help